OpenStudy (anonymous):

Find two consecutive odd numbers such that the sum of one-fifth of the smaller and four-sevenths of the larger is equal to fifty-nine. a) 73 and 75 b) 75 and 77 c) 77 and 79

OpenStudy (anonymous):

let the two no be n and n+2 n is smaller and n+2 is the larger no by the given condiation \[\Large \frac{n}{5}+\frac{4}{7}(n+2)=59\] solve this equation to get value of n

OpenStudy (anonymous):

75 and 77

OpenStudy (anonymous):

yes correct